OpenStudy (anonymous):

WILL MEDAL!! Please help! Find the real or imaginary solutions of each equation using factoring. x^3 + 2x^2+5x+10=0 x^4-10x^2=-9

OpenStudy (irishboy123):

i fancy this one, for sure \[x^4-10x^2=-9\] think: \(z = x^2\)

OpenStudy (xapproachesinfinity):

the first one notice that you can recollect (group) and factor

OpenStudy (xapproachesinfinity):

and the second is as @IrishBoy123 mentioned think of x^2 as another variable z

OpenStudy (xapproachesinfinity):

so in other words you will just get a quadratic equation which is nicer to deal with

OpenStudy (anonymous):

so you're saying do z^2 + 10z

OpenStudy (anonymous):

@xapproachesinfinity ^^

OpenStudy (xapproachesinfinity):

close but not quite \[z^2-10z=-9 \Longrightarrow z^2-10z+9=0 \]

OpenStudy (anonymous):

Ok thank you! Do you know how to do the first one?

OpenStudy (xapproachesinfinity):

so now you are trying to solve \[z^2-10z+9=0\] for z once you found z return back to x

OpenStudy (xapproachesinfinity):

the first is really straight forward \[x^3 + 2x^2+5x+10=0 \Longrightarrow (x^3+2x^2)+(5x+10)=0\] \[x^2(x+2)+5(x+2)=0 \Longrightarrow (x^2+5)(x+2)=0 \]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (xapproachesinfinity):

welcome